# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ Appendix F Iterated Baumslag-Solitar groups
In this appendix we discuss a construction for modifying the Bamsulag-Solitar group ${\sf BS}(n,m)$ in a way which lets us achieve Dehn functions scaling superexponentially in $L$. From the bound (131) this immediately allows us to construct examples of models with nontrivial fragility lengths. Variants of ${\sf BS}(n,m)$ can be constructed whose Dehn functions grow as fast as ${\sf Dehn}(L)\sim\exp^{\lfloor\log_{2}L\rfloor}(1)$ Baumslag (1969), an extremely quickly growing function of $L$. We will however content ourselves with studying the simplest variants, for which the Dehn function scales doubly-exponentially ${\sf Dehn}(L)\sim 2^{2^{L}}$. A family of groups ${\sf BS}^{(2)}(n,m,o,p)$ with this scaling are defined in their simplest presentations by three generators $\mathtt{a},\mathtt{b},\mathtt{c}$, where $\mathtt{a},\mathtt{b}$ satisfy the relations of ${\sf BS}(n,m)$ and $\mathtt{b},\mathtt{c}$ satisfy those of ${\sf BS}(o,p)$:
@@
We will focus throughout on the case $n=o=1,m=p=2$, which is the simplest nontrivial example. For notational brevity, we will write ${\sf BS}^{(2)}(1,2,1,2)$ simply as ${\sf BS}^{(2)}$. The Cayley graph of ${\sf BS}^{(2)}$ consists of an infinite hierarchy of ${\sf BS}$ Cayley graphs. Multiplying by $\mathtt{c}$ increases the “scale” of the ${\sf BS}$ graph generated by $\mathtt{a},\mathtt{b}$, while multiplying by $\mathtt{c}$ decreases the scale. Basic facts about the geometry of ${\mathsf{CG}}_{{\sf BS}^{(2)}}$ such as its growth rate and the size of the identity sector are relatively difficult to address numerically, partly because unlike ${\sf BS}$, ${\sf BS}^{(2)}$ does not admit a simple linear representation, making the evaluation $g(w)$ for arbitrary words $w$ rather involved. [^39] It is in fact an outstanding open question if there exist groups with ${\sf Dehn}(L)>2^{L}$ for which the word problem can be solved in linear time (a matrix representation for $BS^{(2)}$ would answer this question in the affirmative). It is perhaps intuitively reasonable that ${\sf BS}^{(2)}$ should have a superexponential Dehn function, since the $\mathtt{c}$ characters can be used to create exponential expansion of the $\mathtt{b}$ characters, which in turn can create super-exponential expansion of the $\mathtt{a}$ characters. The following result sharpens up this intuition:
@@
This result appears to be well-known and was stated in Ref. Clay and Margalit (2017) without an explicit proof, which we provide below for completeness. The key result we need to complete the proof is known as Britton’s lemma Lyndon et al. (1977), [^40] We thank Tim Riley for suggesting this proof strategy. which is stated as follows:
@@
$G_{\phi}\equiv\langle S,t\,|\,t^{-1}Ht=\phi(H)\rangle.$ (169)
Now any word $w$ on $\{S,t\}^{*}$ can be written in the form
@@
This gives us a way of simplifying any word $w\sim e$ representing the identity in $G_{\phi}$. Case 1 above is trivial. In case 2, we can replace the occurrence of $t^{-1}ht$ with $\phi(h)$, while in case 3 we may replace $tkt^{-1}$ with $\phi^{-1}(k)$. After doing this reduction, we will be left with a new word $w^{\prime}$ which still represents $e$, and can apply the lemma again. This gives us a guarantee that we will always be able to apply a successive series of reductions to eliminate all $t$s from and $w\sim e$ in $\{S,t\}^{*}$ to obtain a word $w_{G}\in\{S\}^{*},w_{G}\sim e$. The application to ${\sf BS}(1,2)$ is clear, where we take $G=\mathbb{Z}$, $H=G$, and $K=2\mathbb{Z}$. We now return to a proof of the Dehn function scaling:
@@
But using our results from our study of ${\sf BS}$, we know that ${\sf Dehn}(\widetilde{w}_{huge}\mathtt{a}\widetilde{w}_{huge}^{-1}\mathtt{a}^{-1})\sim 2^{2^{n}}$. Thus ${\sf Dehn}(L)\geq 2^{2^{L}}$ asymptotically. We now need to provide a matching upper bound. We do this by combining Britton’s lemma with our earlier result about ${\sf BS}$. Note that ${\sf BS}^{(2)}$ can be obtained from ${\sf BS}$ using just the type of extension as appears in Britton’s lemma, where $G={\sf BS},H=\langle\mathtt{b}\rangle,K=\langle\mathtt{b}^{2}\rangle$. Then we know that if we are given $w\sim\mathtt{e}$, $|w|=L$ in ${\sf BS}^{(2)}$, we can obtain a word $w^{\prime}\sim\mathtt{e}$ in ${\sf BS}$ after at most $O(L)$ applications of $\mathtt{c}^{-1}\mathtt{b}\mathtt{c}=\mathtt{b}^{2}$. The maximum amount that $|w|$ can grow by under these substitutions is $O(2^{L})$. Thus an upper bound on ${\sf Dehn}(L)$ in ${\sf BS}^{(2)}$ can be obtained by an upper bound on ${\sf Dehn}(2^{L})$ in ${\sf BS}$. Using our previous result on the latter, we conclude
@@
and thus when combined with the lower bound, we also have that ${\sf Dehn}(L)\lesssim 2^{2^{L}}$. ∎
@@
It is easy generalize the above example to construct groups with even faster growing space and time complexity:
@@
Define the group ${\sf BS}^{(l)}$ through the presentation Clay and Margalit (2017)
@@ Appendix G Detailed analysis of non-group examples
In this Appendix, we provide some detailed analysis studying both the time and space complexity of the non-group examples presented in the main text.
@@
and the entire initial state |$\ket{u}\otimes\ket{v}$ to be in $K_{Q}$ for some large value of $Q$. Let us assume that $v$ is generic enough that a constant fraction of it is 00. We define
@@
where $\mathcal{P}_{n}$ projects onto configurations with $n_{\ast,A}=n$. Let us first suppose that $|A|\ll\log L$. Then, we contend that $p(n,t)>0$ for all $0\leq n\leq|A|$. To show this is simple: in order to allow for $0\leq n\leq|A|$ we need to be able to annihilate all of the parentheses in $A$. Since this can be done in roughly $2^{|A|/2}$ space and $|A|\ll\log L$, configurations with all possibly densities $n_{\ast,A}|$ can be reached. However, when $|A|\gg\log L$ it is no longer that case that all configurations with densities $0\leq n_{\ast,A}\leq|A|$ are achievable. To see this, consider sectors with $Q\gg\exp(|A|)$. Note that in order to have configurations with an amount of charge $n_{\ast,A}=|A|/2+q$, we need to annihilate at least $q$ pairs of parentheses in $A$. If we set the height of the $h$-restriction to be $h=|A|/2-q$ and the value of $Q\gg\exp(|A|)$, there will be at least two contiguous regions. To transport out an amount of charge $2^{q}$ from the contiguous region in $A$ to a contiguous region outside $A$ requires at least $2^{q}/\text{poly}(L)$ space. If $q\gg\log L$ then this is not possible. Therefore, we find that for subsystems of size $\gg\log L$, no configurations with charge $n_{\ast,A}=|A|/2+q$ are reachable with $q\gg\log L$. This implies an unusual property that the reduced density matrix $\rho_{A}$ will not have full rank unless the size of the subsystem is smaller than $\log L$. In this example, the consequences are easily observable since the value of $n_{\ast,A}$ is a local operator.

### 11 Apr 2024 v2

@@ Appendix F Iterated Baumslag-Solitar groups
In this appendix we discuss a construction for modifying the Bamsulag-Solitar group ${\sf BS}(n,m)$ so that the corresponding Dehn function scales superexponentially in $L$. From the spacetime bound (137) this immediately allows us to construct examples of models with superlinear fragility lengths. In general, variants of ${\sf BS}(n,m)$ can be constructed whose Dehn functions grow as fast as ${\sf Dehn}(L)\sim\exp^{\lfloor\log_{2}L\rfloor}(1)$ [55], a rapidly growing function of $L$. We will however study the simplest variants of this class where the Dehn function scales doubly-exponentially ${\sf Dehn}(L)\sim 2^{2^{L}}$. A family of groups ${\sf BS}^{(2)}(n,m,o,p)$ with this scaling are defined in their simplest presentations by three generators $\mathtt{a},\mathtt{b},\mathtt{c}$, where $\mathtt{a},\mathtt{b}$ satisfy the relations of ${\sf BS}(n,m)$ and $\mathtt{b},\mathtt{c}$ satisfy those of ${\sf BS}(o,p)$:
@@
We will focus throughout on the case $n=o=1,m=p=2$, which provides the simplest nontrivial example. For notational brevity, we will write ${\sf BS}^{(2)}(1,2,1,2)$ simply as ${\sf BS}^{(2)}$. The Cayley graph of ${\sf BS}^{(2)}$ consists of an infinite hierarchy of ${\sf BS}$ Cayley graphs. Multiplying by $\mathtt{c}$ (i.e. moving along the $\mathtt{c}$ direction) increases the “scale” of the ${\sf BS}$ Cayley graph generated by $\mathtt{a},\mathtt{b}$, while multiplying by $\mathtt{c}^{-1}$ decreases the scale. Basic facts about the geometry of ${\mathsf{CG}}_{{\sf BS}^{(2)}}$ such as its growth rate and the size of the identity sector are relatively difficult to address numerically, partly because unlike ${\sf BS}$, ${\sf BS}^{(2)}$ does not admit a simple linear representation, making it difficult to compute $g(w)$ for an arbitrary word $w$ [^39] It is in fact an outstanding open question if there exist groups with ${\sf Dehn}(L)>2^{L}$ for which the word problem can be solved in linear time (for instance, a matrix representation for $BS^{(2)}$ would answer this question in the affirmative). One may anticipate that ${\sf BS}^{(2)}$ should have a superexponential Dehn function due to the hierarchical structure of its Cayley graph, since the $\mathtt{c}$ characters can be used to create exponential expansion of the $\mathtt{b}$ characters, which in turn can create doubly-exponential expansion of the $\mathtt{a}$ characters. The following result makes this intuition more precise:
@@
This result appears to be well-known and was stated in Ref. [41] without an explicit proof; we could not find a proof of this statement and thus provide one below for completeness. The key result we need to complete the proof is known as Britton’s lemma [89], [^40] We thank Tim Riley for suggesting this proof strategy. which is stated as follows:
@@
$G_{\phi}\equiv\langle S,t\,|\,t^{-1}Ht=\phi(H)\rangle$ (174)
and wolog express any $w$ on $\{S,t\}^{*}$ in the form
@@
This gives us a way of simplifying any word $w\sim e$ representing the identity in $G_{\phi}$. Case 1 above is trivial. In case 2, we can replace the occurrence of $t^{-1}ht$ with $\phi(h)$, while in case 3 we may replace $tkt^{-1}$ with $\phi^{-1}(k)$. After doing this reduction, we will be left with a new word $w^{\prime}$ which still represents $e$, and can apply the lemma again. Iterating, we are guaranteed to eliminate all $t$s from any $w\sim e$ in $\{S,t\}^{*}$ to obtain a new word $w_{G}\in\{S\}^{*},w_{G}\sim e$. For instance, the group ${\sf BS}(1,2)$ is simply $G_{\phi}$ for the choices $G=\mathbb{Z}$, $H=G$, and $K=2\mathbb{Z}$. We now return to a proof of the Dehn function scaling:
@@
But using our results from our study of ${\sf BS}$, we know that ${\sf Dehn}(\widetilde{w}_{huge}\mathtt{a}\widetilde{w}_{huge}^{-1}\mathtt{a}^{-1})\sim 2^{2^{n}}$. Thus ${\sf Dehn}(L)\geq 2^{2^{L}}$ asymptotically. We now need to provide a matching upper bound. We do this by combining Britton’s lemma with our earlier result about ${\sf BS}$. Note that ${\sf BS}^{(2)}$ can be obtained from ${\sf BS}$ using just the type of extension as appears in Britton’s lemma, where $G={\sf BS},H=\langle\mathtt{b}\rangle,K=\langle\mathtt{b}^{2}\rangle$. Then we know that if we are given $w\sim\mathtt{e}$ where $|w|=L$ in ${\sf BS}^{(2)}$, we can obtain a word $w^{\prime}\sim\mathtt{e}$ in ${\sf BS}$ after at most $O(L)$ applications of $\mathtt{c}^{-1}\mathtt{b}\mathtt{c}=\mathtt{b}^{2}$. The maximum amount that $|w|$ can grow by under these substitutions is $O(2^{L})$. Thus an upper bound on ${\sf Dehn}(L)$ in ${\sf BS}^{(2)}$ can therefore be obtained by an upper bound on ${\sf Dehn}(2^{L})$ in ${\sf BS}$. Using our previous result on the latter, we conclude
@@
and thus when combined with the lower bound, we also have that ${\sf Dehn}(L)\sim 2^{2^{L}}$. ∎
@@
It is easy to generalize the above example to construct groups with even faster growing space and time complexity:
@@
Define the group ${\sf BS}^{(l)}$ through the presentation [41]
@@ Appendix G Detailed analysis of non-group examples
In this appendix we provide some detailed analysis studying both the time and space complexity of the non-group examples presented in the main text.
@@
and the entire initial state |$\ket{u}\otimes\ket{v}$ to be in $K_{Q}$ for some large value of $Q$. Let us assume that $v$ is generic enough that a constant fraction of it is filled with ’00’ characters. We define
@@
where $\mathcal{P}_{n}$ projects onto configurations with $n_{\ast,A}=n$. Let us first suppose that $|A|\ll\log L$. Then, we contend that $p(n,t)>0$ for all $0\leq n\leq|A|$. To show this is simple: in order to allow for $0\leq n\leq|A|$ we need to be able to annihilate all of the parentheses in $A$. Since this can be done in roughly $2^{|A|/2}$ space and $|A|\ll\log L$, configurations with all possibly densities $n_{\ast,A}|$ can be reached. However, when $|A|\gg\log L$ it is no longer that case that all configurations with densities $0\leq n_{\ast,A}\leq|A|$ are achievable. To see this, consider sectors with $Q\gg\exp(|A|)$. Note that in order to have configurations with $n_{\ast,A}=|A|/2+q$, we need to annihilate at least $q$ pairs of parentheses in $A$. If we set the height of the $h$-restriction to be $h=|A|/2-q$ and the value of $Q\gg\exp(|A|)$, there will be at least two contiguous regions. To transport out an amount of charge $2^{q}$ from the contiguous region in $A$ to a contiguous region outside $A$ requires at least $2^{q}/\text{poly}(L)$ space. If $q\gg\log L$ then this is not possible. Therefore, we find that for subsystems of size $\gg\log L$, no configurations with charge $n_{\ast,A}=|A|/2+q$ are reachable with $q\gg\log L$. This implies an unusual property that the reduced density matrix $\rho_{A}$ will not have full rank unless the size of the subsystem is smaller than $\log L$. In this example, the consequences are easily observable since the value of $n_{\ast,A}$ is a local operator.